Nite Labs <Chapters 1. 2, and 3 Problem Quiz Problem 259 17 a Review 1 Constants I Periodic TablePart A In major league baseball, the pitchers mound is 60 feet from the batter tValue Units
Answer #1

1 feet =0.3048 m

distance between pitcher and batter

d=60 ft =60*0.3048 =18.288 m

1 mph =0.44704 m/s

Speed

V=85*0.44704 =38 m/s

Therefore time taken by the ball to reach the batter

t=d/V =18.288/38

t=0.4813 s =0.48 s